OpenStudy (ballislife):

Write an equation in slope-intercept form of the line that passes through (−5, 19) and(5, 13).

OpenStudy (gokuporter):

First find the slope which is \[\frac{ -6 }{ 10 }\]

OpenStudy (ballislife):

I just can't figure out how to find the y-intercept

OpenStudy (the_fizicx99):

y - y1 = m(x-x1) y - 19 = -3/5(x-(-5)) y - 19 = -3/5(x +5) y - 19 = -3/5x - 3 +19 +19 y = -3/5x + 16